OpenStudy (anonymous):

a business finds the average cost required to complete a project is $416 with a standard deviation of $57. a sample of 36 project cost reports was selected at random. assuming that the distribution of the costs is normal, what is the probability that the average for this sample is between $400 and $425?

OpenStudy (anonymous):

z1=(400-416)/57=-.28 z2=(425-416)/57=.16 now find P(z2)-P(z1)

OpenStudy (anonymous):

Would you happen to know the formula for that? manually or in excel?

OpenStudy (anonymous):

i got .1739 does that sound about right?